OpenStudy (anonymous):

A pendulum with length of 1.4 meters swings through an angle of 30 degree .How far does the bob at the end of the pendulum travel as goes from left to right?

OpenStudy (anonymous):

i'm trying to remember for length of an arc given radius and angle.. length = r * c where c = angle in radians - yup so l = 1.4 * 30pi/ 180 = 1.4 * pi/6